A hockey player makes a slap shot, exerting a

force of 37.2 N on the hockey puck for 0.147 s.

What impulse is given to the puck?

Answer in units of kgm/s.


1
Expert's answer
2022-03-01T11:12:59-0500

The impulse

"\Delta p=F\Delta \\
=37.2*0.147=4.8\:\rm kg\cdot m/s"